# What We Talk About When We Talk About LMs: Implicit Paradigm Shifts and the Ship of Language Models

### v1

@@ 2 Related Work
Our work provides a complement to these ongoing threads. As discussed, we raise a novel scenario about the transitions within a lasting concept (Ship of LMs), which to our knowledge has not been explored. We examine the use of such terms as LMs, providing quantitative interpretations of how (and how much) our beliefs and common grounds have evolved. In some sense, our work can be also seen as a meta-analysis of the various works studying certain elements (e.g. “the era of LLM”, “stages of statistical Machine Translation”, “ChatGPT’s impact”, etc.) We integrate these valuable findings to highlight a new question about the procedures: how exactly did we forge these of key elements into practice and eventually to our norms of language?
@@ 3.2 Retrieving the Mentions of LMs
To investigate the Ship of LMs problem, we start by extracting and analyzing keywords and relevant entities, a common backbone method for analysis Hou et al. (2019); Pramanick et al. (2023). For a sample paper to be related to LMs, the writing could utilize two types of mentions: (1) the collective concept of “language models”, implying the context as a generalizable discussion, and (2) the names of specific models, indicating what models are exactly considered in a limited scope.
@@ 3.2.1 Notations
since “language model” is the substring of most of its subcategories, e.g., “large language models”, “Korean language models”, or “language modeling”, and searching for “language model” covers all such variations. We also include the most common acronyms, “LLM” (Large Language Models) and “PLM” (Pre-trained Language Models). The construction of $\mathcal{M}$ is elaborated in §3.2.2.
@@ 3.2.2 Constructing $\mathcal{M}$ from the text
To compose the final list, we sorted names following a simple heuristic (illustrated in Figure 1):
  1. 1.
    When we encounter a new model $m$ not in $\mathcal{M}$, we add $m$ to $\mathcal{M}$ and initialize its alias list as $[m]$;
  2. 2.
    Additional aliases of an entry $m_{j}$ are appended to its list;
  3. 3.
    Variations of existing entries are recorded but not added as an alias;
  4. 4.
    Candidates that are not the name of a model (e.g., BLEU) are discarded.
For each entry of $m$, we also manually retrieved the original paper or documentation to determine if there is an explicit dependency on another model. In all, $\mathcal{M}$ has a total of 98 model entries and 146 aliases. With the two keyword lists, $\mathcal{L}$ and $\mathcal{M}$, we are ready to examine how LMs are resolved and extract diachronic patterns.
Our dataset and code will be open-sourced, and we provide the full list of models involved and the LLM prompts in Appendix A.1 and A.2.
@@ 4 Experiments and Findings
LMs have been steadily gaining more attention from the field. Zhao et al. (2023) reports that papers containing the key phrase “Language Model” have increased from less than 400 pre-2019 to around 10,000 in 2023. We observe a similar trend with a finer-grained search and the focus on main conference publications in the NLP domain (Figure 2(a)). At ACL 2020, 35% of the papers contain at least one LM mention from $\mathcal{L}$ (we refer to this portion as LM-related papers). Since then, this proportion has had a smooth, continuous growth of approximately $5\%$ (additive) per conference, hitting 84% just three years later at EMNLP 2023.
@@ 4.1 Wind in the Sails: Surging Mentions, Speeding Conclusions
We consider the average $N^{\mathcal{L}}$ of all papers at a conference ($\bar{N^{\mathcal{L}}}$). If the case above is true, $\bar{N^{\mathcal{L}}}$ would remain generally unchanged in both the LM-related and non-related groups. Thus, the ratio of LM-related papers determines $\bar{N^{\mathcal{L}}}$, and we can draw an estimate from the scale of the first data point ($\bar{N^{\mathcal{L}}}=4.29$ for $35\%$ at ACL 2020). For instance, 54% of papers at EMNLP 2021 are LM-related, which is 1.54×1.54$\times$1.54 × that of ACL 2020 (35%). We scale $\bar{N^{\mathcal{L}}}$ with the same ratio, $1.54\times 4.29\approx 6.56$, as the estimate for EMNLP 2021.
Fig. 2(b) compares the actual $\bar{N^{\mathcal{L}}}$ and the value estimated in this way. This turns out to be a surprisingly good fit for the first half of the data. Within the first 5 conferences, the deviations between estimated and actual value are consistently less than 10% and often close to 0, e.g., we estimated $6.56$ for EMNLP 2021 where the actual value is $6.59$. For this period, LMs gained more attention as a topic in this period, but language describing this term remained similar. Metaphorically, the composition of the Ship remains the same, but it has more wind in its sails.
@@
First, we see evidence supporting our prior observations on the patterns of $N^{\mathcal{L}}$. Earlier conferences form a cluster where no significant difference is noted; yet, starting 2022, every conference has a significantly higher ${N^{\mathcal{L}}}$ than most or even all of its predecessors. However, the distributions of $N$ tell a distinct story. For most pairs, there is little or no evidence for a difference in distribution. There also isn’t a similar line that divides the earlier and most recent conferences. For instance, the distribution of $N$ for EMNLP in 2023 does not have a significant difference with that in 2020, despite all its specialties. In fact, we even see an opposite case: conferences in 2022 and 2023 – the exact time of the super-linear boosts of $\bar{N^{\mathcal{L}}}$ – contain significantly fewer model mentions than before. In other words, we arrive at the conclusions faster: the information conveyed via specific models has not increased, but more is drawn about LMs collectively. Possible reasons include more focus on the whole concept of the Ship rather than maintaining its parts, or on specific planks than the wholesale renovation of the vessel; more efforts would be needed to understand the exact cause.
@@ 4.2 Oak, Pine, or Cedar Planks: Which models are we talking about?
With the exploding usage of the term LMs comes wider variation in the use cases and context around them. To go deeper, we must consider what writers refer to when they include LM in a paper (i.e., what the Ship is like at a certain point and how do they [re]build the ship, and not just whether it sails.) Based on all individual $N_{m}$ and the hierarchy of components, we obtain the exact number of the appearances of each model by matching their aliases in the text. Thus, we put together the collective compositions for each conference and visualize them as Sunburst charts, where the component sizes correspond to their share. We show a representative comparison of EMNLP 2020 and 2023 in Figure 4, and display full results in Appendix B.
@@
How much remains as the replacement of earlier components goes on? We calculate Jaccard similarity to quantify how much is shared between the composition of conferences, shown in Figure 5. We observe that Jaccard similarity between conferences monotonically decreases for subsequent conferences, which matches the Ship of LMs case where constituent parts change over time. For two consecutive conferences, the Jaccard similarity is usually only 71% to 85%; the index quickly drops to 45% to 57% with an interval of just two years, and to 24% to 31% in three years. We also note that dissimilarity is rapidly increasing, with EMNLP 2023 sharing a 52% Jaccard similarity with ACL 2023 just half a year ago, 42% with EMNLP in 2022, and $\leq$38with all other predecessors. With the representing models thoroughly reshuffled in as short as $<$5 years, the “shelf life” of our conclusions and knowledge of LM has seen a new low, thus bringing unprecedented challenges for long-term tasks and literature studies.
@@
Figure 6 displays the proportion of publications with absolute majority components for the full data (left) and the top 25% with the highest $N^{\mathcal{L}}$ (right). The most notable components are marked with the same color as in Fig. 4. Other models are shown collectively as the grey bar (“others”), and the proportion with no single majority model is denoted with a striped pattern. We observe that around 80% of papers contain an absolute majority model. Specifically, we get a glance at the astounding traction of BERT before more recent paradigm shifts: it dominated up to 60.5% of all papers and 68.6% of the most LM-centered ones.
@@ 5 Concluding Remarks
In this work, we go over the past and present of the enduring term of Language Model(s), based on an original dataset from the latest major conferences. We sort out the subtle, continuous shifts in the practical meaning of LMs, and witness how the retrofits eventually accumulate to a brand new Ship of Language Models. We checked up the accelerated transmutation of our use of the term beyond the “routine updates” of the Ship, whereas actual referents did not sync with the pace. We quantify and visualize the drastic change in the planks and timber, emphasizing the shortened period of reconstruction and the presence of dominant components. Finally, we highlight the snowballing context difference between the LM-centered research and the more peripheral applications and its consequences.
Our work seeks to delineate the shape of the drifting concepts in the tide of time; yet, snapshots at the galas still do not seem to catch up with the differentials as much originally as a decade’s but now within months or weeks. We have in fact seen more interesting signs, standing at another epochal crossroads of the ChatGPT era: the unprecedented dissimilarity of compositions, or on the other hand, a more diverse, multipolar representation in model choices, to name a few. Perhaps, what is crafted from our hand signals the Age of Sail that we are yet to know – where it’d be too much for an epic Argo from the fanta-seas to explore.
@@
Our findings depict the most dramatic reshuffles of models, while the nuance of LMs is not limited to the model level. The scope can extend to the broader, underlying architecture level, as readers might have noticed the dominance of the Transformer models. Conversely, it is also intriguing to zoom in to the variations of models, as models are eventually instantiated by specific versions just like LMs are resolved to models.
The understanding of scientific progress can also benefit from various other genres. Causal links between our work and the identified key factors in prior works could draw a more vivid picture of the field’s trajectories. Qualitative studies, such as interviews with the paper authors, can also be an informative complement that delves deeper into our beliefs and explores where the patterns in different groups come from. Aside from the LM-specific discussions, we also highlight the broad existence of the Ship of Theseus scenario in various other terms, topics, and fields.
@@
While our methodologies can be naturally applied to similar scenarios, we would like to note that the current analyses and implementations have limitations. First, although we have gone over various findings in recent years with the most dramatic advances of LMs, a holistic overview of a research topic, let alone an entire field, is not covered by the scope of a few years. It would be meaningful if these most rapid changes could be connected with the decades of conceptualization and exploration preceding the engineering breakthroughs. Similarly, using main conference papers at the major international venues as a proxy of the NLP community has its limitations. We encourage future works to take broader consideration of the essential contributions that are less represented by the relatively convergent selections of such venues, e.g., regional conferences on dialect or indigenous language.
@@ Appendix A The dataset construction pipeline
We provide more details about the semi-automated dataset creation process: the full list of models involved, and the prompts used for the automated part of model name identification. The full dataset and more details will be published for future use in the final version of this project.
A.1 Full list of models
ChatGPT, GPT-3, GPT-4, BERT, T5, GPT-3.5, GPT-2, LLaMA, RoBERTa, PaLM, CLIP, BART, XLM-R, Alpaca, BLOOM, mT5, InstructGPT, mBERT, GPT-J, Flan-T5, OPT, Codex, COMET, ELECTRA, Longformer, mBART, SimCSE, BLOOMZ, BigBird, BLIP, DeBERTa, CodeT5, Switch Transformer, Vicuna, T0, PEGASUS, LSTM, ALBERT, DPR, Macaw, LXMERT, SpanBERT, TinyBERT, ViLBERT, TransE, RotatE, XLM, Linformer, kNN-LM, kNN-MT, REALM, RETRO, GraphCodeBERT, Sentence-BERT, RNN, HyperCLOVA, CodeGen, Dolly, Pythia, LaMDA, FLAN, BLIP-2, XLNet, GPT, ELMo, BioBERT, DialoGPT, RemBERT, PaLM 2, DistilBERT, SciBERT, ClinicalBERT, M2M100, GloVe, LASER, word2vec, fastText, LaBSE, CNN, wav2vec, UNITER, MASS, MT-DNN, BlenderBot, DistMult, OFA, CMLM, HRED, ERNIE, ConveRT, MiniLM, Galactica, RuleTakers, Claude, LayoutLM, ST-DNN, IRNet
A.2 Prompts for automated model name extraction

### v2

@@ 2 Related Work
Our work complements these ongoing threads. As discussed, we raise a novel scenario about the transitions within a lasting concept (Ship of LMs), which to our knowledge has not been explored. We examine the use of such terms as LMs, providing quantitative interpretations of how (and how much) our beliefs and common grounds have evolved. In some sense, our work can also be seen as a meta-analysis of the various works studying certain elements (e.g. “the era of LLM”, “stages of statistical Machine Translation”, “ChatGPT’s impact”, etc.) We integrate these valuable findings to highlight a new question about the procedures: how exactly did we forge these of key elements into practice and eventually into our norms of language?
@@ 3.2 Retrieving the Mentions of LMs
To investigate the Ship of LMs problem, we start by extracting and analyzing relevant keywords and entities, a common backbone method for analysis Hou et al. (2019); Pramanick et al. (2023). For a sample paper to be related to LMs, the writing could utilize two types of mentions: (1) the collective concept of “language models”, implying the context as a generalizable discussion, and (2) the names of specific models [^3] We consider the broadest definition and scope of LMs; see Appendix A.1 for a more detailed note., indicating what models are exactly considered in a limited scope.
@@ 3.2.1 Notations
since “language model” is the substring of most of its subcategories, e.g., “large language models”, “Korean language models”, or “language modeling”, and searching for “language model” [^4] Here we include other upper-case forms, e.g., “Language Models”. This is implemented similar to Aliases in §3.2.2. covers all such variations. We also include the most common acronyms, “LLM” (Large Language Models) and “PLM” (Pre-trained Language Models). The construction of $\mathcal{M}$ is elaborated in §3.2.2.
@@ 3.2.2 Constructing $\mathcal{M}$ from the text
To compose the final list, we sorted names following a simple heuristic (illustrated in Appendix A.2). When we encounter a new model $m$ not in $\mathcal{M}$, we add $m$ to $\mathcal{M}$ and initialize its alias list as $[m]$. Additional aliases of an entry $m_{j}$ are appended to its list. Variations of existing entries are recorded but not added as an alias, and candidates that are not the name of a model (e.g., BLEU) are discarded.
For each entry of $m$, we also manually retrieved the original paper or documentation to determine if there is an explicit dependency on another model. In all, $\mathcal{M}$ has a total of 97 model entries and 146 aliases. With the two keyword lists, $\mathcal{L}$ and $\mathcal{M}$, we are ready to examine how LMs are resolved and extract diachronic patterns.
Our dataset and code will be open-sourced, and we provide the full list of models involved and the LLM prompts in Appendix A.3 and A.4.
@@ 4 Experiments and Findings
LMs have been steadily gaining more attention from the field. Zhao et al. (2023) reports that papers containing the key phrase “Language Model” have increased from less than 400 pre-2019 to around 10,000 in 2023. We observe a similar trend with a finer-grained search in the NLP domain (Figure 1(a)). At ACL 2020, 35% of the papers contain at least one LM mention from $\mathcal{L}$ (we refer to this portion as LM-related papers). Since then, this proportion has had a smooth, continuous growth of approximately $5\%$ (additive) per conference, hitting 84% just three years later at EMNLP 2023.
@@ 4.1 Wind in the Sails: Surging Mentions, Speeding Conclusions
We consider the average $N^{\mathcal{L}}$ of all papers at a certain conference, written as $\bar{N^{\mathcal{L}}}$. With the division of LM-related and non-related papers, $\bar{N^{\mathcal{L}}}$ can be calculated as the (weighted) average of the two groups. If the above assumption holds true, $\bar{N^{\mathcal{L}}}$ can be estimated as follows: On the one hand, non-LM-related papers always have $N^{\mathcal{L}}=0$ by definition. On the other hand, within the LM-related group, the average of $N^{\mathcal{L}}$ would also be stable across conferences since the use of the term LMs remains generally unchanged as hypothesized. As both groups – constituting the full set of papers at a conference when combined – see no qualitative change in their within-group $N^{\mathcal{L}}$, the determining factor for the overall average $\bar{N^{\mathcal{L}}}$ is the ratio of the two groups’ size (as the weight), i.e., the percentage of LM-related papers. Thus, we can draw an estimate from the scale of the first data point ($\bar{N^{\mathcal{L}}}=4.29$ for $35\%$ LM-related at ACL 2020). For instance, 54% of papers at EMNLP 2021 are LM-related, which is 1.54×1.54$\times$1.54 × that of ACL 2020 (35%). We can thus scale $\bar{N^{\mathcal{L}}}$ with the same ratio, $1.54\times 4.29\approx 6.56$, as the estimate for EMNLP 2021.
Fig. 1(b) compares the actual $\bar{N^{\mathcal{L}}}$ and the value estimated in this way. This turns out to be a surprisingly good fit for the first half of the data. Within the first 5 conferences, the deviations between estimated and actual value are consistently less than 10% and often close to 0. For this period, LMs gained more attention as a topic in this period, but language describing this term remained similar. Metaphorically, the composition of the Ship remains the same, but it has more wind in its sails.
@@
First, we see evidence supporting our prior observations on the patterns of $N^{\mathcal{L}}$. Earlier conferences form a cluster where no significant difference is noted; yet, starting 2022, every conference has a significantly higher ${N^{\mathcal{L}}}$ than most or even all of its predecessors. However, the distributions of $N$ tell a distinct story. For most pairs, there is little or no evidence for a difference in distribution. There also isn’t a similar line that divides the earlier and most recent conferences. For instance, the distribution of $N$ for EMNLP in 2023 does not have a significant difference with that in 2020, despite all its specialties. In fact, we even see an opposite case: conferences in 2022 and 2023 – the exact time of the super-linear boosts of $\bar{N^{\mathcal{L}}}$ – contain significantly fewer model mentions than before. In other words, we arrive at the conclusions faster: the information conveyed via specific models has not increased, but more is drawn about LMs collectively. Various factors might account for this trend: It might be that we are increasingly proficient in maintaining the Ship, or that more weight is put on the holistic concept of the Ship compared with closer examinations on the level of specific models, etc. More effort would be needed to discern the exact cause, and deeper understandings of the local context of LMs (e.g. similar to Jurgens et al., 2018) would be especially relevant for future work.
@@ 4.2 Oak, Pine, or Cedar Planks: Which models are we talking about?
With the exploding usage of the LM terms comes wider variation in the use cases and context around them. To go deeper, we must consider what writers refer to when they include LM in a paper (i.e., quantify what the Ship is like and how it’s being updated at a certain point, and not just whether it sails.) Based on all individual $N_{m}$ and the hierarchy of components, we obtain the exact number of the appearances of each model by matching their aliases in the text. Thus, we put together the collective compositions for each conference and visualize them as Sunburst charts, where the component sizes correspond to their share. We show a representative comparison of EMNLP 2020 and 2023 in Figure 3, and display full results in Appendix B.
@@
How much remains as the replacement of earlier components goes on? We calculate the Jaccard similarity between compositions of conferences to quantify how much is shared across any two conferences, shown in Figure 4. We observe that Jaccard similarity between conferences monotonically decreases for subsequent conferences, which matches the Ship of LMs case where its parts change over time. For two consecutive conferences, the Jaccard similarity is usually only 71% to 85%; the index quickly drops to 45% to 57% with an interval of just two years, and to 24% to 31% in three years. We also note that dissimilarity is rapidly increasing, with EMNLP 2023 sharing a 52% Jaccard similarity with ACL 2023 just half a year ago, 42% with EMNLP in 2022, and $\leq$38with all other predecessors. With the representing models thoroughly reshuffled in as short as $<$5 years, the “shelf life” of our conclusions and knowledge of LM has seen a new low, thus bringing unprecedented challenges for long-term tasks and literature studies.
@@
Figure 5 displays the proportion of publications with absolute majority components for the full data (left) and the top 25% with the highest $N^{\mathcal{L}}$ (right). The most notable components are marked with the same color as in Fig. 3. Other models are collected as the grey bar (“others”), and the proportion with no absolute majority is denoted with a striped pattern. Around 80% of the papers contain an absolute majority model. Specifically, we get a glance at the astounding traction of BERT before more recent paradigm shifts: it dominated up to 60.5% of all papers and 68.6% of the most LM-centered ones.
@@ 5 Concluding Remarks
In this work, we go over the past and present of the enduring term of Language Model(s), based on an original dataset from the latest major conferences. We sort out the subtle, continuous shifts in the practical meaning of LMs, and witness how the retrofits eventually accumulate to a brand new Ship of Language Models and the nuances in the actual referents. We quantify and visualize the drastic change in the planks and timber, emphasizing the shortened period of reconstruction and the presence of dominant components. Finally, we highlight the snowballing context difference between the LM-centered research and the more peripheral applications and the consequences of that.
Approaching the most recent epochal crossroads where revolutionary shifts can happen within months or weeks, we have in fact seen more interesting signs: the unprecedented dissimilarity of compositions, or on the other hand, a more diverse, multipolar representation in model choices, to name a few. Perhaps what is crafted from our hands has even brought up throught the Age of Sail, where a now steam-engined Argo is ready to set sail towards the exciting yet mysterious fanta-seas.
@@
Our findings depict the most dramatic reshuffles of models, while the nuances of LMs can be found similarly at the underlying architecture level, or different variations of models. Besides, causal links between our work and the identified key factors in prior works may help depict the field’s trajectories, and qualitative studies such as interviews with paper authors can also help understand the different patterns and beliefs. Beyond the LM-specific discussions, we also highlight the broad existence of the Ship of Theseus scenario in various other terms, topics, and fields.
@@
While our methodologies can be naturally extended, we would like to note that the current analyses and implementations have limitations. First, although we have gone over various findings in recent years with the most dramatic advances of LMs, a holistic overview of a research topic, let alone an entire field, is not covered by the scope of a few years. It would be meaningful if these most rapid changes could be connected with the decades of conceptualization and exploration preceding the engineering breakthroughs. Similarly, using main conference papers at the major international venues as a proxy of the NLP community has its limitations. We encourage future works to take broader consideration of the essential contributions that are less represented by the relatively convergent selections of such venues, e.g., regional conferences on dialect or indigenous language.
@@ Appendix A The dataset construction pipeline
We provide more details about the semi-automated dataset creation process, including the illustration of the workflow, the full list of models involved, and the prompts used for the automated part of model name identification. The full dataset and more details will be published for future use in the final version of this project.
A.1 A note on the scope of “LMs”
In the rigid sense, LMs shall correspond directly to the task of next-word prediction by enlisting the probabilities of each possible next-word (adapted from the prose of Jurafsky and Martin (2023)). However, practical referents and contexts of “language modeling” in the real world have been broad and complex. In fact, in Jurafsky and Martin’s work cited above, the definition of LMs (given as a direct statement) has already been broader than previous senses in the prose:
Models that assign probabilities to upcoming words, or sequences of words in general, are called language models or LMs.
To accommodate the underspecification and the even more vague real-world usages, we thus consider the “broadest” definitions that are historically involved. This would include text embedding schemes that have essentially been used in the place of “LMs”, key parts of an LM-like system where the system is not otherwise named, or artifacts created in the “LM” way but are not exactly a “model”. (Note that this work concerns “what we talk about when we talk about LMs”, not just “which LM we talk about when we talk about LMs”.)
This is especially worth noting since the terminology of “LM” is not always the most prevalent. As an important instance, “word embeddings/vectors” had once been a heated topic; it is both true that (1) the terms are not equivalent to “LMs” strictly speaking, and that (2) it was a very common (if not default) practice to build (e.g. add a linear layer) on top of the embeddings to obtain probabilities/likelihood of the text in question. This is essentially close to “LM” as defined, though we rarely referred to the entire system this way but would have often used wordings like “word vectors + [classification/prediction/…]” back then. To this end, we do consider these “broader” senses because they have long been used for the same sense and in the same context. In fact, our motivations align with such situations: the communities’ description and perception of a concept might have undergone many changes and variations, but we are yet to know what exactly has changed, to what extent, and how varied it is.
We exemplify our considerations below with some keywords in the list:
  • Word2vec is the algorithm where the training goal is exactly to assign a max likelihood/probability of the current text (and not to forget N-gram LMs is indeed a type of LM too – there goes the Ship of LMs!);
  • The reasons for other word embedding schemes like GloVE has been addressed by the discussions so far;
  • RNN, LSTM, etc. are the neural networks that generate continuous text representations, and they have been prominently applied for the exact task of probabilistic next-word/sentence generation.
  • There are also individual considerations. For instance, CNN, CLIP, and wav2vec have been under discussion regarding the multimodality issues, and we eventually decided to keep (potentially) multimodal keywords.
A.2 Classifying potential model names
See Figure 8.
A.3 Full list of models
ChatGPT, GPT-3, GPT-4, BERT, T5, GPT-3.5, GPT-2, LLaMA, RoBERTa, PaLM, CLIP, BART, XLM-R, Alpaca, BLOOM, mT5, InstructGPT, mBERT, GPT-J, Flan-T5, OPT, Codex, COMET, ELECTRA, Longformer, mBART, SimCSE, BLOOMZ, BigBird, BLIP, DeBERTa, CodeT5, Switch Transformer, Vicuna, T0, PEGASUS, LSTM, ALBERT, DPR, Macaw, LXMERT, SpanBERT, TinyBERT, ViLBERT, XLM, Linformer, kNN-LM, kNN-MT, REALM, RETRO, GraphCodeBERT, Sentence-BERT, RNN, HyperCLOVA, CodeGen, Dolly, Pythia, LaMDA, FLAN, BLIP-2, XLNet, GPT, ELMo, BioBERT, DialoGPT, RemBERT, PaLM 2, DistilBERT, SciBERT, ClinicalBERT, M2M100, GloVe, LASER, word2vec, fastText, LaBSE, CNN, wav2vec, UNITER, MASS, MT-DNN, BlenderBot, DistMult, OFA, CMLM, HRED, ERNIE, ConveRT, MiniLM, Galactica, RuleTakers, Claude, LayoutLM, ST-DNN, IRNet
A.4 Prompts for automated model name extraction